OpenStudy (lena772):

A system of equations is shown below. -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (lena772):

@ranga can you please help me?

OpenStudy (ranga):

-3x + 7y = -16 ---- (1) -9x + 5y = 16 ---- (2) "Replace one equation with the sum of that equation and a multiple of the other." Let us replace (2) by adding (2) to two times (1): 2 times (1): -6x + 14y = -32 Add (2): -9x + 5y = 16 -------------------- -15x + 19y = -16 So the second pair of equations are: -3x + 7y = -16 ---- (1) -15x + 19y = -16 ----- (3) For B) Solve (1) and (2) for x,y Solve (1) and (3) for x,y Show they are the same solutions.

OpenStudy (lena772):

(1) (2) (3) ??? what do these mean??

OpenStudy (lena772):

@ranga

OpenStudy (lena772):

Sorry I get it now! Thank you so so much!!!